# Gas Transfer Between the Inner 3-kpc Disk and the Galactic Central Molecular Zone

### v1

@@ Abstract
We uncovered a more tilted molecular gas structure with highly negative velocities located near the dust lane. Our observations show that the approaching gas flows from the overshoot process are captured by the bar gravitational and then flows towards the Galactic Central Molecular Zone (CMZ) through the bar channel. The recycling gas from the overshoot effect, in conjunction with freshly accreted gas from the inner 3-kpc disk, accumulates significantly near and regions by adopting a bar length of 3.2–3.4 kpc. Importantly, within these regions, there are frequent collisions and substantial angular momentum exchanges between gas flows with different trajectories. In this scenario, the DISSIPATION processes arising from interactions between colliding flows, together with the varying torques induced by the non-axisymmetric bar, effectively transfer the angular momentum of viscous gas outward, thereby driving the molecular gas to settle into the CMZ within about 3 orbital periods. A long-term gas inflow with an average rate of , coupled with intense transient accretion events that exceed the average rate by several times due to the overshoot effect, significantly regulates the gas distribution, physical properties, and dynamical evolution of the CMZ. These new findings provide robust observational evidence for elucidating the intricate dynamics of molecular gas flows towards the CMZ. Our observations suggest that gas dynamics have a significant impact on the secular evolution of both the Milky Way and the extragalactic gas-rich galaxies.
@@ 1 Introduction
The Central Molecular Zone (CMZ), which may originate from the gas inflow driven by the Galactic bar, hosts a huge amount of molecular gas (i.e., 2–6, Dahmen et al., 1998; Launhardt et al., 2002; Ferrière et al., 2007; Battersby et al., 2024) within 200–300 pc of Sgr A∗. As the closest observable galactic nucleus, the CMZ provides us with a wealth of critical information about gas cycling in large-scale Galactic dynamics, star formation in extreme environments, and the various physical processes related to Galactic evolution (see the comprehensive reviews of Bryant & Krabbe, 2021; Henshaw et al., 2023). Such studies on the Galactic CMZ are also of great significance for understanding the gas dynamics in other galaxies (e.g., the case NGC 1365 in Schinnerer et al., 2023), as well as the evolution of galaxies and their own nuclei (e.g., Leroy et al., 2021; Stuber et al., 2023; Schinnerer & Leroy, 2024).
@@
Molecular gas is the dominant gas component in the inner region of the Milky Way (e.g., Koda et al., 2016). Through spectral observations of molecular gas, we can investigate the distribution, kinematics, and other properties of molecular clouds (MCs) in detail. Although 12CO does not reveal the internal density structure of MCs well, it is a good indicator of the extended structure of MCs and a useful tracer of the total mass of molecular gas. Additionally, the simultaneous observation of the 13CO and C18O lines can provide useful chemical information on MCs. By combining multi-wavelength data from molecular gas and other tracers of stars, we are well placed to investigate the star formation in the Milky Way. So far, millimeter and sub-millimeter CO observations have been particularly important in studying molecular gas characteristics towards the Galactic CMZ.
@@
Benefiting from the high quality 12CO and 13CO data from the Milky Way Imaging Scroll Painting project (MWISP, Su et al., 2019), Su et al. (2024) have identified many CO structures with unique morphological and kinematic features distributed along the major axis of the Galactic bar (see the Galactic stellar bar in, e.g., Wegg et al., 2015; Bland-Hawthorn & Gerhard, 2016; Simion et al., 2017; Shen & Zheng, 2020; Lucey et al., 2023; Vislosky et al., 2024). The bar channel traced by the CO dust lane connects the Galactic disk at the Galactocentric distance of 3.2–3.4 kpc and the CMZ at kpc. On the other hand, the length of the bar dust lanes is generally consistent with the radial scale length of the Galactic exponential disk (i.e., and kpc, see Figure 13 in Su et al., 2021), likely indicating that the bar dynamics have a significant impact on the properties of the molecular gas in the inner Galaxy (e.g., distribution, morphology, kinematics, chemistry, and evolution). Our study provides solid evidence that the large-scale molecular gas inflow is driven by the Galactic bar towards the CMZ (see additional observations and simulations, e.g., Fux, 1999; Rodriguez-Fernandez & Combes, 2008; Sormani et al., 2015; Sormani & Barnes, 2019).
@@
The CO flows in the bar channel, driven by the non-axisymmetric potential, become a secular and stable gas reservoir that feeds star-formation in the CMZ and potentially supplies the supermassive black hole (SMBH) in the Galactic center (GC). The gas inflow rate onto the CMZ is estimated to be about yr-1 in a transport period of 24 Myr by adopting an average inflow velocity of for the large-scale gas lane (Su et al., 2024). According to the MWISP CO data, the pattern speed of supports that the Milky Way is currently in a slow-bar scenario, which is consistent with recent studies (e.g., Hilmi et al., 2020; Clarke & Gerhard, 2022; Wheeler et al., 2022; Zhang et al., 2024a).
@@
Despite much progress has been made in multi-wavelength observations and numerical simulations, some open issues still need to be carefully investigated so that we can gain a more accurate understanding of the gas dynamics in the inner region of the Milky Way. How does the gas with significant angular momentum transfer from the Galactic disk several kpc away to the CMZ and even the GC? How the molecular gas fuels that subsequently form stars are affected by the Galactic large-scale dynamics? What is the relationship between the gas inflow rate and the efficiency of the star formation in the CMZ? And what role do massive stellar winds, photo-ionization, radiation pressure, and supernova feedback play in gas accumulation of the CMZ?
@@
Our observations are instantaneous and freeze-frame. Can the observations accurately reveal the dynamical evolution of molecular gas flows towards the CMZ over time? In some specific studies, numerical simulations often simplify and discard many of the details of the large-scale gas flows in which the CMZ is embedded. Can some of the striking and interesting features in the simulations be verified by observations? In the paper, we investigate the dynamical nature of gas flows towards the CMZ based on the MWISP CO survey. Thanks to the large-scale MWISP data with high spatial resolution and sensitivity, we are able to capture many spatial and kinematic details of the molecular gas flow towards the CMZ, which have never been revealed in previous studies.
@@
In Section 2, we show that the tilted overshoot gas flows from the far dust lane converge in the near bar channel, and eventually settle into the CMZ. In Section 3, a schematic diagram is constructed to illustrate both the overall dynamic behavior and detailed gas flow patterns towards the CMZ. We suggest that the dissipation processes resulting from collisions between large-scale gas flows are critical for removing the gas’s angular momentum and facilitating its inward migration by considering the overshoot effect. In Section 4, we propose that the non-axisymmetric gravitational potential of the bar play a significant role in regulating the gas dynamics within the 3 kpc of the Galaxy during its secular evolution. Finally, we give a summary in Section 5.
@@ 2 Overshoot gas from the Far Dust Lane
In Figure 1, the kinematic properties of the large-scale CO structures are well delineated towards the CMZ in the region of and . Besides the CO inflow driven by the Galactic bar (hereafter DL1; see the details in Su et al., 2024), two additional large-scale CO structures have been revealed flowing towards the CMZ along the near dust lane (i.e., DL2 and DL3; also refer to Liszt, 2006, 2008; Sormani et al., 2019). These gas flows are receding relative to the observer, so we identify them with red arrows in the - diagrams.
@@
Unexpectedly, a new CO structure with the most negative velocity is discovered below the Galactic plane according to the large-scale MWISP data (the blue long-arrow in the lower panel of Figure 1). The new kinematic CO structure consists of many filamentary MCs and exhibits a gradient that increases with Galactic longitude. Obviously, the CO feature is not related to the near 3 kpc arm (see the blue long-line in the lower panel of Figure 1; also refer to Dame & Thaddeus, 2008; Reid et al., 2016, 2019; Kumar et al., 2025). On the contrary, the large-scale gas structure coexists with DL1 in a similar - range. More precisely, the approaching flow is roughly parallel to or slightly below the near dust lane (i.e., see the colors map for the approaching gas and the white contours for the receding DL1 gas in Figure 2c). We define the large-scale approaching CO gas as the overshoot flows from the far dust lane (e.g., Liszt, 2006; Wallace et al., 2022).
@@
In Figure 2c, an intriguing filamentary cloud associated with the overshoot gas flow is found to be located in the G7 region () with a large velocity gradient of by adopting a distance of 6 kpc (see the parameters of the geometry in Su et al., 2024). The value is at least two times larger than the mean velocity gradient of the near dust lane (i.e., for DL1), indicating rapid deceleration of overshoot gas in the G7 region. The feature stems from the fact that the overshoot gas is undergoing strong tidal processes when it moves in the Galactic bar potential. As molecular gas flows pass through the apocentre of elongated elliptical orbits, MCs experience extremely strong torques and shear forces, resulting in significant geometric deformation from intense tidal effects (e.g., Kruijssen et al., 2019). Moreover, a considerable amount of overshoot gas is being pulled back by the Galactic bar potential and re-enters the bar channel towards the CMZ, which is well delineated by the blue short-arrows (OG1 and OG2) and the red arrows (DL2 and DL3) in Figure 1.
@@
The overshoot flows are changing their trajectories in response to the Galactic bar potential, i.e., (1) from OG1 below the Galactic plane to DL2 above the plane (the upper panel of Figure 1) and (2) from OG2 at to DL3 at (the lower panel of Figure 1). Indeed, we distinctly illustrates that the large-scale multilayer shock structures with varying velocities are located in coherent but slightly different positions, indicating that the motion state of a series of overshoot flows with similar trajectories has changed dramatically in the vicinity of this region (see the shock structures indicated in blue, red, and green colors close to the G5 region at and ; Figure 2a). The data show that the overshoot gas is re-accelerating from to and even more at about the half of the Galactic bar (i.e., at the middle of the near dust lane DL1 of or 1.7–1.5 kpc) when it passes through the bar channel. The observations appear to be broadly consistent with the scenario of significant gas accumulation around the inner Lindblad resonance (ILR; e.g., 1.6 kpc in Sormani et al., 2024).
@@
Collisions between gas flows are also revealed by observations. Combined with complementary distributions between the high-velocity gas (i.e., white contours for the DL2 inflows in Figure 2b) and the low-velocity overshoot gas (colors for the OG3 flows; also see Figures 6 and 7 in Eden et al., 2020), for example, the prominent CO protrusions evidently demonstrate that the two gas flows in opposite directions are colliding violently at the G3 region of (, ). Interestingly, the receding gas of DL3 passing below G3 (see red color near –, – in Figure 2b) also appears to collide with the surrounding overshoot gas, causing G3 to exhibit a deformed structure at (, ). The observations are not only clear, but also align well with the cloud-cloud collision scenarios discussed in literatures (e.g., see Furukawa et al., 2009; Inoue & Fukui, 2013; Gong et al., 2017; Enokiya et al., 2021; Fukui et al., 2021; Gramze et al., 2023). According to the large-scale bow shock perpendicular to the dust lane in Figure 2a, the G5 region at 1.6 kpc is another excellent example of the collision between gas flowing along the bar channel (i.e., white contours for high-velocity receding gas in DL1) and the approaching overshoot flows passing through the bar channel (i.e., the blue and green colors for the low- and middle-velocity gas in OG2+DL3; also see Akhter et al., 2021). We believe that the approaching gas flows from overshoot are bombarding the large-scale receding gas streams in the vicinity of the near dust lane.
@@
It is noteworthy that the large-scale shock structure near G5 (indicated by the yellow dotted line in Figure 2a) spans approximately 2.7 degrees on the sky plane, corresponding to a projected physical size of about 320 pc at the distance of 6.8 kpc (see the geometry of the bar lanes in Su et al., 2024). Here, the size of 320 pc may provide a reliable measurement of the cross-sectional diameter of the bar channel, which is comparable to the size of the CMZ (e.g., see Henshaw et al., 2023). Indeed, the shock features (i.e., velocity jump) on the bar region can be well reproduced in simulations (e.g., Athanassoula, 1992; Kim & Stone, 2012; Feng et al., 2024). And our observations can provide additional constrains for hydrodynamic simulations of the Milky Way case. We propose that the strength of the inflowing gas, as well as gravitational potential (e.g., bar strength, axis ratio, and central mass distribution) and local ISM properties (e.g., turbulence level, magnetic field strength, and star formation activity), influence the size and other properties of the Galactic CMZ.
@@
At least two small breaks along DL1 can be identified in the - diagram (see the tips of OG1 and OG2 in the lower panel of Figure 1). The features can naturally be explained by the convergence of the overshoot gas from the far dust lane to the DL1 fresh gas from the 3–4 kpc region, leading discernible changes in the velocity gradient and gas distribution of the whole gas inflows along the near dust lane. The scenario also agrees with the presence in which the overshoot gas with lower velocities is mainly concentrated upstream and the DL1 gas with higher velocities is enhanced downstream (i.e., the slight displacement between the overshoot gas represented by the colors and the DL1 gas by white contours in Figure 2c). Moreover, the gradual velocity-shift features are indeed discerned along and/or near the OG1 and OG2 in the - diagrams (Figure 1). For the first time, our observations provide detailed evidence that large amounts of overshoot gas from the far dust lane is captured by the Galactic bar potential at these different and then inflowing towards the CMZ along or parallel to the bar channel (e.g., DL2 and DL3).
@@
According to the Gaussian decomposition and clustering algorithm (e.g., see Henshaw et al., 2019; Riener et al., 2019; Zhang et al., 2024b), we identified the MCs toward the CMZ by using the smoothing parameters of 1=1.8 and 2=5.4 for the MWISP 12CO data with the 3D grid of . It is straightforward to pick out the DL1 samples from the near dust lane and the overshoot samples from the far dust lane based on the - diagrams (Figure 1). Accordingly, the intermediate samples are selected from MCs with between the DL1 and the overshoot samples, i.e., MCs in DL2, DL3, G3, G5, and near the head part of OG1, OG2, and OG3, excluding those from the blended regions. Table 1 displays the statistical properties of three types of MCs with angular sizes larger than 50 arcmin2.
@@
By employing the method consistent with our earlier work (Su et al., 2024), we consistently derive the CO-to-H2 conversion factor of cm-2(K km s. The value is about half of the commonly used value (e.g., Dame et al., 2001; Strong et al., 2004; Bolatto et al., 2013; Kohno & Sofue, 2024) and is close to (K km s for the Galactic bulge region (e.g., Schultheis et al., 2014). Due to shocks driven by the bar and frequent cloud-cloud collisions, there may be multiple gas components with different temperatures and densities near the bar channel (e.g., Nilipour et al., 2024), resulting in systematic changes in the molecular gas state. Generally, in the strong shock and collision regions of G3 and G5, MCs with wide line broadenings often have a relatively small ratio between 13CO and 12CO intensity, which is about half or less of the value of the normal MCs in the Galactic plane (e.g., CO)/CO)0.2, see Tokuyama et al., 2019; Wang et al., 2023). The same is true of statistics for other gas flows associated with the Galactic bar (i.e., the large-scale MC samples with less contamination in the near dust DL1 and overshoot flows; see Table 1). Based on observations, the statistical result of the smaller CO)/CO) likely represents the lower in the complex shock environments of the inner Galaxy.
@@
The Milky Way represents an excellent case study for understanding gas cycling in galaxy evolution. Detailed observations have provided a wealth of data and information about the Milky Way, such as the 3D spatial distribution of gas flows, the precise geometric structure of the bar, and other relevant key physical parameters (e.g., the gas inflow rate, the outflow rate, and the star formation rate in the CMZ, etc.). These detailed observational insights are crucial for constructing and constraining theoretical models, as they significantly advance our understanding of gas dynamics in galaxies and their secular evolution. Intriguingly, external galaxies also exhibit similar overshooting features (NGC 1097 and NGC 1365; see e.g., Sormani et al., 2023; Whitmore et al., 2023; Schinnerer et al., 2023). Exploring whether gas cycling processes and associated physical parameters (e.g., geometry of bar and overshooting gas, distribution of gravitational potential, and gas inflow rate, etc.) in external galaxies and comparing them with those in the Milky Way can elucidate the fundamental role of gas cycling in the broader context of galaxy evolution.
@@
In summary, the molecular gas is a good indicator of the gravitational potential of the Galactic bar because the gas flows can be easily and quickly affected by the change in the gravitational torque of the rotating bar. The dynamical processes dominated by the gravitational potential of the bar cause the overshoot gas to decelerate, accumulate, and collide violently at certain locations, systematically changing the gas motion and its overall distribution towards the CMZ. We propose that gas dynamics play a crucial role in studying several key issues, such as the redistribution of inflowing material and the regulation of overall gas inflow efficiency (Section 3), as well as the subsequent star formation efficiency in the CMZ (Section 4).
3 Overall Dynamics of the Gas Flows towards the CMZ

### v2

# Gas Transfer Between the Inner 3 kpc Disk and the Galactic Central Molecular Zone

@@ Abstract
We uncovered a more tilted molecular gas structure with highly negative velocities located near the dust lane. Our observations show that the approaching gas flows from the overshoot process are captured by the gravitational potential of the bar and then flow toward the Galactic central molecular zone (CMZ) through the bar channel. The recycled gas from the overshoot effect, in conjunction with freshly accreted gas from the inner 3 kpc disk, accumulates significantly near and regions by adopting a bar length of 3.2–3.4 kpc. Importantly, within these regions, there are frequent collisions and substantial angular momentum exchanges between gas flows with different trajectories. In this scenario, the DISSIPATION processes arising from interactions between colliding flows, together with the varying torques induced by the nonaxisymmetric bar, effectively transfer the angular momentum of viscous gas outward, thereby driving the molecular gas to settle into the CMZ within about three orbital periods. A long-term gas inflow with an average rate of , coupled with intense transient accretion events that exceed the average rate by several times due to the overshoot effect, significantly regulates the gas distribution, physical properties, and dynamical evolution of the CMZ. These new findings provide robust observational evidence for elucidating the intricate dynamics of molecular gas flows toward the CMZ. Our observations suggest that gas dynamics have a significant impact on the secular evolution of both the Milky Way and the extragalactic gas-rich galaxies.
@@ 1 Introduction
The central molecular zone (CMZ), which may originate from the gas inflow driven by the Galactic bar, hosts a huge amount of molecular gas (i.e., 2–6; Dahmen et al., 1998; Launhardt et al., 2002; Ferrière et al., 2007; Battersby et al., 2024) within 200–300 pc of Sgr A∗. As the closest observable galactic nucleus, the CMZ provides us with a wealth of critical information about gas cycling in large-scale Galactic dynamics, star formation in extreme environments, and the various physical processes related to Galactic evolution (see the comprehensive reviews of Bryant & Krabbe, 2021; Henshaw et al., 2023). Such studies on the Galactic CMZ are also of great significance for understanding the gas dynamics in other galaxies (e.g., the case NGC 1365 in Schinnerer et al., 2023), as well as the evolution of galaxies and their own nuclei (e.g., Leroy et al., 2021; Stuber et al., 2023; Schinnerer & Leroy, 2024).
@@
Molecular gas is the dominant gas component in the inner region of the Milky Way (e.g., Koda et al., 2016). Through spectral observations of molecular gas, we can investigate the distribution, kinematics, and other properties of molecular clouds (MCs) in detail. Although 12CO does not reveal the internal density structure of MCs well, it is a good indicator of the extended structure of MCs and a useful tracer of the total mass of molecular gas. Additionally, the simultaneous observation of the 13CO and C18O lines can provide useful chemical information on MCs. By combining multiwavelength data from molecular gas and other tracers of stars, we are well placed to investigate the star formation in the Milky Way. So far, millimeter and submillimeter CO observations have been particularly important in studying molecular gas characteristics toward the Galactic CMZ.
@@
Benefiting from the high-quality 12CO and 13CO data from the Milky Way Imaging Scroll Painting (MWISP) project (Su et al., 2019), Su et al. (2024) have identified many CO structures with unique morphological and kinematic features distributed along the major axis of the Galactic bar (see the Galactic stellar bar in, e.g., Wegg et al., 2015; Bland-Hawthorn & Gerhard, 2016; Simion et al., 2017; Shen & Zheng, 2020; Lucey et al., 2023; Vislosky et al., 2024). The bar channel traced by the CO dust lane connects the Galactic disk at the Galactocentric distance of 3.2–3.4 kpc and the CMZ at kpc. On the other hand, the length of the bar dust lanes is generally consistent with the radial scale length of the Galactic exponential disk (i.e., and kpc, see Figure 13 in Su et al., 2021), likely indicating that the bar dynamics have a significant impact on the properties of the molecular gas in the inner Galaxy (e.g., distribution, morphology, kinematics, chemistry, and evolution). Our study provides solid evidence that the large-scale molecular gas inflow is driven by the Galactic bar toward the CMZ (see additional observations and simulations; e.g., Fux, 1999; Rodriguez-Fernandez & Combes, 2008; Sormani et al., 2015; Sormani & Barnes, 2019).
@@
The CO flows in the bar channel, driven by the nonaxisymmetric potential, become a secular and stable gas reservoir that feeds star-formation in the CMZ and potentially supplies the supermassive black hole (SMBH) in the Galactic center (GC). The gas inflow rate onto the CMZ is estimated to be about yr-1 in a transport period of 24 Myr by adopting an average inflow velocity of for the large-scale gas lane (Su et al., 2024). According to the MWISP CO data, the pattern speed of supports that the Milky Way is currently in a slow-bar scenario, which is consistent with recent studies (e.g., Hilmi et al., 2020; Clarke & Gerhard, 2022; Wheeler et al., 2022; Zhang et al., 2024a).
@@
Despite much progress being made in multiwavelength observations and numerical simulations, some open issues still need to be carefully investigated so that we can gain a more accurate understanding of the gas dynamics in the inner region of the Milky Way. How does the gas with significant angular momentum transfer from the Galactic disk several kiloparsecs away to the CMZ and even the GC? How are the molecular gas fuels that subsequently form stars affected by the Galactic large-scale dynamics? What is the relationship between the gas inflow rate and the efficiency of the star formation in the CMZ? And what role do massive stellar winds, photoionization, radiation pressure, and supernova feedback play in the gas accumulation of the CMZ?
@@
Our observations are instantaneous and freeze-frame. Can the observations accurately reveal the dynamical evolution of molecular gas flows toward the CMZ over time? In some specific studies, numerical simulations often simplify and discard many of the details of the large-scale gas flows in which the CMZ is embedded. Can some of the striking and interesting features in the simulations be verified by observations? In this paper, we investigate the dynamical nature of gas flows toward the CMZ based on the MWISP CO survey. Thanks to the large-scale MWISP data with high spatial resolution and sensitivity, we are able to capture many spatial and kinematic details of the molecular gas flow toward the CMZ that have never been revealed in previous studies.
@@
In Section 2, we show that the tilted overshoot gas flows from the far dust lane converge in the near bar channel and eventually settle into the CMZ. In Section 3, a schematic diagram is constructed to illustrate both the overall dynamic behavior and detailed gas flow patterns toward the CMZ. We suggest that the dissipation processes resulting from collisions between large-scale gas flows are critical for removing the gas’s angular momentum and facilitating its inward migration by considering the overshoot effect. In Section 4, we propose that the nonaxisymmetric gravitational potential of the bar plays a significant role in regulating the gas dynamics within 3 kpc of the Galaxy during its secular evolution. Finally, we give a summary in Section 5.
@@ 2 Overshoot Gas from the Far Dust Lane
In Figure 1, the kinematic properties of the large-scale CO structures are well delineated toward the CMZ in the region of and . Besides the CO inflow driven by the Galactic bar (hereafter DL1; see the details in Su et al., 2024), two additional large-scale CO structures have been revealed flowing toward the CMZ along the near dust lane (i.e., DL2 and DL3; also refer to Liszt, 2006, 2008; Sormani et al., 2019). These gas flows are receding relative to the observer, so we identify them with red arrows in the – diagrams.
@@
Unexpectedly, a new CO structure with the most negative velocity is discovered below the Galactic plane according to the large-scale MWISP data (the long blue arrow in the lower panel of Figure 1). The new kinematic CO structure consists of many filamentary MCs and exhibits a gradient that increases with Galactic longitude. Obviously, the CO feature is not related to the near 3 kpc arm (see the long blue line in the lower panel of Figure 1; also refer to Dame & Thaddeus, 2008; Reid et al., 2016, 2019; Kumar et al., 2025). On the contrary, the large-scale gas structure coexists with DL1 in a similar – range. More precisely, the approaching flow is roughly parallel to or slightly below the near dust lane (i.e., see the colors for the approaching gas and the white contours for the receding DL1 gas in Figure 2(c)). We define the large-scale approaching CO gas as the overshoot flows from the far dust lane (e.g., Liszt, 2006; Wallace et al., 2022).
@@
In Figure 2(c), an intriguing filamentary cloud associated with the overshoot gas flow is found to be located in the G7 region () with a large velocity gradient of by adopting a distance of 6 kpc (see the parameters of the geometry in Su et al., 2024). The value is at least 2 times larger than the mean velocity gradient of the near dust lane (i.e., for DL1), indicating rapid deceleration of overshoot gas in the G7 region. The feature stems from the fact that the overshoot gas is undergoing strong tidal processes when it moves in the Galactic bar potential. As molecular gas flows pass through the apocenter of elongated elliptical orbits, MCs experience extremely strong torques and shear forces, resulting in significant geometric deformation from intense tidal effects (e.g., Kruijssen et al., 2019). Moreover, a considerable amount of overshoot gas is being pulled back by the Galactic bar potential and reenters the bar channel toward the CMZ, which is well delineated by the short blue arrows (OG1 and OG2) and the red arrows (DL2 and DL3) in Figure 1.
@@
The overshoot flows are changing their trajectories in response to the Galactic bar potential, i.e., (1) from OG1 below the Galactic plane to DL2 above the plane (the upper panel of Figure 1) and (2) from OG2 at to DL3 at (the lower panel of Figure 1). Indeed, we distinctly illustrate that the large-scale multilayer shock structures with varying velocities are located in coherent but slightly different positions, indicating that the motion state of a series of overshoot flows with similar trajectories has changed dramatically in the vicinity of this region (see the shock structures indicated in blue, red, and green colors close to the G5 region at and ; Figure 2(a)). The data show that the overshoot gas is reaccelerating from to and even more at about half of the Galactic bar (i.e., at the middle of the near dust lane DL1 of or 1.7–1.5 kpc) when it passes through the bar channel. The observations appear to be broadly consistent with the scenario of significant gas accumulation around the inner Lindblad resonance (ILR; e.g., 1.6 kpc in Sormani et al., 2024).
@@
Collisions between gas flows are also revealed by observations. Combined with complementary distributions between the high-velocity gas (i.e., white contours for the DL2 inflows in Figure 2(b)) and the low-velocity overshoot gas (colors for the OG3 flows; also see Figures 6 and 7 in Eden et al., 2020), for example, the prominent CO protrusions evidently demonstrate that the two gas flows in opposite directions are colliding violently at the G3 region of (, ). Interestingly, the receding gas of DL3 passing below G3 (see red color near –, – in Figure 2(b)) also appears to collide with the surrounding overshoot gas, causing G3 to exhibit a deformed structure at (, ). The observations are not only clear but also align well with the cloud–cloud collision scenarios discussed in the literature (e.g., see Furukawa et al., 2009; Inoue & Fukui, 2013; Gong et al., 2017; Enokiya et al., 2021; Fukui et al., 2021; Williams et al., 2022; Gramze et al., 2023). According to the large-scale bow shock perpendicular to the dust lane in Figure 2(a), the G5 region at 1.6 kpc is another excellent example of the collision between gas flowing along the bar channel (i.e., white contours for high-velocity receding gas in DL1) and the approaching overshoot flows passing through the bar channel (i.e., the blue and green colors for the low- and middle-velocity gas in OG2+DL3; also see Akhter et al., 2021). We believe that the approaching gas flows from overshoot are bombarding the large-scale receding gas streams in the vicinity of the near dust lane.
@@
It is noteworthy that the large-scale shock structure near G5 (indicated by the yellow dotted line in Figure 2(a)) spans approximately 27 on the sky plane, corresponding to a projected physical size of about 320 pc at the distance of 6.8 kpc (see the geometry of the bar lanes in Su et al., 2024). Here, the size of 320 pc may provide a reliable measurement of the cross-sectional diameter of the bar channel, which is comparable to the size of the CMZ (e.g., see Henshaw et al., 2023). Indeed, the shock features (i.e., velocity jump) on the bar region can be well reproduced in simulations (e.g., Athanassoula, 1992; Kim & Stone, 2012; Feng et al., 2024). And our observations can provide additional constraints for hydrodynamic simulations of the Milky Way case. Potential signs of star formation are also observed near the apex boundary of the large-scale shock structure (see GMC M4.7-0.8 in Butterfield et al., 2025). This result may support the hypothesis that as the gas flow approaches the orbital apocenter (this work), the decrease in orbital velocity induces gas compression (density enhancement), thereby triggering localized star formation (Chaves-Velasquez et al., 2025). We propose that the strength of the inflowing gas, as well as gravitational potential (e.g., bar strength, axis ratio, and central mass distribution) and local interstellar medium properties (e.g., turbulence level, magnetic field strength, and star formation activity), influence the size and other properties of the Galactic CMZ.
@@
At least two small breaks along DL1 can be identified in the – diagram (see the tips of OG1 and OG2 in the lower panel of Figure 1). The features can naturally be explained by the convergence of the overshoot gas from the far dust lane to the DL1 fresh gas from the 3–4 kpc region, leading discernible changes in the velocity gradient and gas distribution of the whole gas inflows along the near dust lane. The scenario is also consistent with the observational fact that lower-velocity overshooting gas is mainly concentrated upstream, while the-higher-velocity DL1 gas becomes enhanced downstream (i.e., the slight displacement between the overshoot gas represented by the colors and the DL1 gas by white contours in Figure 2(c)). Moreover, the gradual velocity-shift features are indeed discerned along and/or near OG1 and OG2 in the – diagrams (Figure 1). For the first time, our observations provide detailed evidence that large amounts of overshoot gas from the far dust lane are captured by the Galactic bar potential at these different and then inflow toward the CMZ along or parallel to the bar channel (e.g., DL2 and DL3).
@@
According to the Gaussian decomposition and clustering algorithm (e.g., see Henshaw et al., 2019; Riener et al., 2019; Zhang et al., 2024b), we identified the MCs toward the CMZ by using the smoothing parameters of 1=1.8 and 2=5.4 for the MWISP 12CO data with the 3D grid of . It is straightforward to pick out the DL1 samples from the near dust lane and the overshoot samples from the far dust lane based on the - diagrams (Figure 1). Accordingly, the intermediate samples are selected from MCs with between the DL1 and the overshoot samples, i.e., MCs in DL2, DL3, G3, and G5 and near the head part of OG1, OG2, and OG3, excluding those from the blended regions. Table 1 displays the statistical properties of three types of MCs with angular sizes larger than 50 arcmin2.
@@
By employing the method consistent with our earlier work (Su et al., 2024), we consistently derive the CO-to-H2 conversion factor of cm-2(K km s. The value is about half of the commonly used value (e.g., Dame et al., 2001; Strong et al., 2004; Bolatto et al., 2013; Kohno & Sofue, 2024) and close to (K km s for the Galactic bulge region (e.g., Schultheis et al., 2014). Due to shocks driven by the bar and frequent cloud–cloud collisions, there may be multiple gas components with different temperatures and densities near the bar channel (e.g., Nilipour et al., 2024), resulting in systematic changes in the molecular gas state. Generally, in the strong shock and collision regions of G3 and G5, MCs with wide line broadenings often have a relatively small ratio between 13CO and 12CO intensity, which is about half or less of the value of the normal MCs in the Galactic plane (e.g., CO)/CO)0.2; see Tokuyama et al., 2019; Wang et al., 2023). The same is true of statistics for other gas flows associated with the Galactic bar (i.e., the large-scale MC samples with less contamination in the near dust DL1 and overshoot flows; see Table 1). Based on observations, the statistical result of the smaller CO)/CO) likely represents the lower in the complex shock environments of the inner Galaxy.
@@
The Milky Way represents an excellent case study for understanding gas cycling in galaxy evolution. Detailed observations have provided a wealth of data and information about the Milky Way, such as the 3D spatial distribution of gas flows, the precise geometric structure of the bar, and other relevant key physical parameters (e.g., the gas inflow rate, the outflow rate, the star formation rate in the CMZ, etc.). These detailed observational insights are crucial for constructing and constraining theoretical models, as they significantly advance our understanding of gas dynamics in galaxies and their secular evolution. Intriguingly, external galaxies also exhibit similar overshooting features (NGC 1097 and NGC 1365; see, e.g., Sormani et al., 2023; Whitmore et al., 2023; Schinnerer et al., 2023). Exploring the gas cycling processes and associated physical parameters (e.g., geometry of the bar and overshooting gas, distribution of gravitational potential, gas inflow rate, etc.) in external galaxies and comparing them with those in the Milky Way can elucidate the fundamental role of gas cycling in the broader context of galaxy evolution.
@@
In summary, the molecular gas is a good indicator of the gravitational potential of the Galactic bar because the gas flows can be easily and quickly affected by the change in the gravitational torque of the rotating bar. The dynamical processes dominated by the gravitational potential of the bar cause the overshoot gas to decelerate, accumulate, and collide violently at certain locations, systematically changing the gas motion and its overall distribution toward the CMZ. We propose that gas dynamics play a crucial role in studying several key issues, such as the redistribution of inflowing material and the regulation of overall gas inflow efficiency (Section 3), as well as the subsequent star formation efficiency in the CMZ (Section 4).
3 Overall Dynamics of the Gas Flows toward the CMZ